# Using Synchronic Definitions and Semantic Relations to Classify Semantic Change Types

### v1

@@ 2.3 Classification in Historical Semantics
The lack of resources tailored to train models on these intricate semantic change has prevented computational work thus far. Semantic change often entail a complete loss of previous meanings, making most existing lexical resources insufficient, as they typically miss the historical meanings. Although datasets exist for phenomena like metaphors, metonymies, or analogies, they often concentrate on the novel use of word meanings and omit conventional examples Maudslay and Teufel (2022) thus limiting the models’ ability to predict semantic change types. While there exists no resources for training models, in this paper, we release a first benchmark for testing models on their ability to detect semantic change types given definitions.
@@ 7.1 Semantic Relatedness and Homonymy
Figure 3: Ratio of human judgments for each class.
For each usage in the English portion of the SemEval-2020 Task 1, we generate a definition. We then predict word sense relationships using the definitions for all the usage pairs for which human judgments are available. The word sense relationships include homonymy, co-hyponymous transfer (co-hyponym), generalization (hyperonym), and specialization (hyponym). The bar graph in Figure 3 represents, for each predicted class, the distribution of the existing human judgments based on the semantic relatedness scale [^8] It is important to note that the type classification model used in the study is not designed to identify identical meanings, which corresponds to a human judgment of 4..
@@ Acknowledgments
This work has in part been funded by the project Towards Computational Lexical Semantic Change Detection supported by the Swedish Research Council (2019–2022; contract 2018-01184), and in part by the research program Change is Key! supported by Riksbankens Jubileumsfond (under reference number M21-0021). The computational resources were provided by the National Academic Infrastructure for Supercomputing in Sweden (NAISS), partially funded by the Swedish Research Council through grant agreement no. 2022-06725.
@@ Appendix A LSC-CTD Benchmark
In Table 6, we present an excerpt from the LSC-CTD Benchmark, a manually curated dataset of dictionary-like definition pairs for words that experienced semantic change. It includes Blank’s original annotations detailing the target word followed by a language tag (lt: Latin, vlt: Vulgar Latin) and German glosses of the original, primary meaning- (Old Meaning) and of the derived, secondary meaning (New Meaning). We translate Blank’s glosses into English using Google Translate’s API [^9] https://cloud.google.com/ (between double quotation marks). Following this, we use the prompt shown in Figure 7 to generate definitions that are more akin to those found in a dictionary, employing ChatGPT [^10] We use the gpt-4 API chat.openai.com (between single quotation marks). This prompt includes some examples we initially set up, with the actual generation occurring through few-shot learning. Furthermore Blank’s dataset contains the alleged cause of the semantic change and, not shown here, the higher-level association relation that exists at the conceptual level. Finally, a linguist with expertise in historical linguistics and semantics manually reviewed and refined the collection of generated translations and definitions, replacing or modifying them as necessary.
@@ Appendix C WN Statistics
In Table 5 we report the statistics we use to train the sense relationship model. The table includes for each split (train, dev, and test set) the unique number of synsets, the number of pairs of definitions, and the number of pairs of definitions for each class (hyponym, hypernym, co-hyponym, homonym, and antonym). Additionally, the number of distinct pairs for each part of speech (nouns, verbs, adjectives, and adverbs) is reported.

### v3

@@ 2.3 Classification in Historical Semantics
The lack of resources tailored to train models on these complex semantic changes has prevented computational work thus far. Semantic change often entails a complete loss of previous meanings, making most contemporary computational lexical resources insufficient, as they typically miss the historical meanings. Although datasets exist for phenomena like metaphors, metonymies, or analogies, they often concentrate on the novel use of word meanings and omit conventional examples Maudslay and Teufel (2022) thus limiting the models’ ability to predict semantic change types. While there exists no resources for training models, in this paper, we release a first benchmark for testing models on their ability to detect semantic change types given definitions.
@@ 7 Evaluation on Semantic Change
Figure 3: Ratio of human judgments for each class.
@@ 7.1 Semantic Relatedness and Homonymy
For each usage in the English portion of the SemEval-2020 Task 1, we generate a definition. We then predict word sense relationships using the definitions for all the usage pairs for which human judgments are available. The word sense relationships include homonymy, co-hyponymy (co-hyponymous transfer) , hyperonymy (generalization), hyponymy (specialization) and antonymy (auto-antonymy). The bar graph in Figure 3 represents, for each predicted class, the distribution of the existing human judgments based on the semantic relatedness scale [^8] It is important to note that the type classification model used in the study is not designed to identify identical meanings, which corresponds to a human judgment of 4..
@@ Acknowledgments
This work has in part been funded by the project Towards Computational Lexical Semantic Change Detection supported by the Swedish Research Council (2019–2022; contract 2018-01184), and in part by the research program Change is Key! supported by Riksbankens Jubileumsfond (under reference number M21-0021) and by an FWO junior postdoctoral grant (1281222N). The computational resources were provided by the National Academic Infrastructure for Supercomputing in Sweden (NAISS), partially funded by the Swedish Research Council through grant agreement no. 2022-06725.
@@ Appendix A LSC-CTD Benchmark
In Table 6, we present an excerpt from the LSC-CTD Benchmark, a manually curated dataset of dictionary-like definition pairs for words that experienced semantic change. It includes Blank’s original annotations detailing the target word followed by a language tag (lt: Latin, vlt: Vulgar Latin) and German glosses of the original, primary meaning- (Old Meaning) and of the derived, secondary meaning (New Meaning). We translate Blank’s glosses into English using Google Translate’s API (between double quotation marks). Following this, we use the prompt shown in Figure 7 to generate definitions that are more akin to those found in a dictionary, employing gpt-4 API (between single quotation marks). This prompt includes some examples we initially set up, with the actual generation occurring through few-shot learning. Furthermore Blank’s dataset contains the alleged cause of the semantic change and, not shown here, the higher-level association relation that exists at the conceptual level. Finally, a linguist with expertise in historical linguistics and semantics manually reviewed and refined the collection of generated translations and definitions, replacing or modifying them as necessary.
@@ Appendix C WN Statistics
In Table 5 we report the statistics we use to train the sense relationship model. The table includes for each split (train, dev, and test set) the unique number of synsets, the number of pairs of definitions, and the number of pairs of definitions for each class (hyponym, hyperonym, co-hyponym, homonym, and antonym). Additionally, the number of distinct pairs for each part of speech (nouns, verbs, adjectives, and adverbs) is reported.